What mass of Gallium chloride would react with excess fluorine to produce 78.9 g of chlorine? Gallium fluoride is the other product.
1
Expert's answer
2020-05-24T02:36:49-0400

GaCl3 + 3/2 F2 = GaF3 + 3/2 Cl2

78.9 g of Cl2 is equal to 78.9 g / 71 g/mole = 1.111 mole of chlorine

number of moles of GaCl3 is 2/3 of number of moles of Cl2, i.e. 2/3*1.111 = 0.741 mole of GaCl3.

Mass of gallium chloride is 0.741 mole * 176.08 g/mole = 130.45 g